OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation 3x^2 + 6y^2 = 18 Center: (0, 0); Vertices:(0,-sqrt6) (0,sqrt6) ; Foci: (0,-sqrt3) (0,sqrt3) Center: (0, 0); Vertices: (-sqrt6, 0) (sqrt6, 0) ; Foci: (-sqrt3,0) (sqrt3,0) Center: (0, 0); Vertices: (-6, 0), (6, 0); Foci: (-3sqrt3, 0) (3sqrt3, 0) Center: (0, 0); Vertices: (0, -6), (0, 6); Foci: (0,-3sqrt3) (0,3sqrt3)

OpenStudy (kropot72):

The first step is to put the equation in standard form \[\frac{x ^{2}}{a ^{2}}+\frac{y ^{2}}{b ^{2}}=1\] To do this, divide both sides of the equation by 18 and simplify. Can you do that?

OpenStudy (anonymous):

1/6 + 1/3 =1

OpenStudy (anonymous):

actually x^2/6 + y^2/3 =1 ??

OpenStudy (kropot72):

Good work! Looking at the standard form of the equation for an ellipse again \[\frac{x ^{2}}{a ^{2}}+\frac{y ^{2}}{^{b ^{2}}}=1\] the vertices are at \[(\pm a, 0)\] Can you find the vertices now?

OpenStudy (kropot72):

Hint:\[a ^{2}=6\] \[a=\sqrt{6}\]

OpenStudy (kropot72):

The vertices are at (-a, 0) and (a, 0). Just plug in sqrt6 where there is a.

OpenStudy (kropot72):

@n14r96 Are you there?

OpenStudy (anonymous):

\[3x ^{2}+6y ^{2}=18\] \[\frac{x ^{2} }{6 }+\frac{y ^{2} }{3 }=1\] \[centre is \left( 0,0 \right)\] \[vertices are \left( -\sqrt{6},0 \right) and \left( \sqrt{6},0 \right) ,\left(0- \sqrt{3} \right),\left(0,\sqrt{3} \right)\] \[compare with \frac{ x ^{2} }{ a ^{2} }+\frac{y ^{2} }{ b ^{2} }=1,b ^{2}=a ^{2}\left( 1-e ^{2} \right)\] \[3=6\left( 1-e ^{2} \right),\frac{ 3 }{ 6 }=1-e ^{2},\frac{ 1 }{ 2 }-1=-e ^{2},e=\frac{ 1 }{ \sqrt{2} }\] Foci are (-ae,0) and (ae,0) \[\left( \frac{- \sqrt{6} }{\sqrt{2} },0 \right)and \left( \frac{ \sqrt{6} }{\sqrt{2} },0 \right)\] \[\left( -\sqrt{3},0 \right)and \left( \sqrt{3},0 \right)\]